OpenStudy (anonymous):

Draw a right triangle. Roll a die two times and record the numbers you roll. On the drawing of your right triangle, label the legs of the triangle using the numbers you rolled. Use the Pythagorean Theorem to find the length of the hypotenuse and write out the steps you performed using words. You must show all work to receive credit. Here's an example below. Notice the dice have values above the values you will have on your six-sided dice. Please use your own original values from rolling your dice:

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@zepdrix u got this man cmon

zepdrix (zepdrix):

bam step 1 done.

zepdrix (zepdrix):

bust out them dice man

OpenStudy (anonymous):

i got that much. say i roll a 3 and a 5

zepdrix (zepdrix):



OpenStudy (anonymous):

okay. now i have no idea what to do. this is where i get stumped

zepdrix (zepdrix):

Pythagorean Theorem states that the sum of the square of the legs is equal to the square of the hypotenuse,\[\large\rm 3^2+5^2=hypotenuse^2\]

zepdrix (zepdrix):

Maybe we should call the hypotenuse by a letter, so it's a little easier to work with,\[\large\rm 3^2+5^2=x^2\]We need to solve for x.

OpenStudy (anonymous):

so \[9+25=x^2\]